OpenStudy (anonymous):

Hi! What is the equation in point-slope form of the line passing through (1, 2) and (2, 5)? a. (y - 2) = -3(x - 1) b. (y - 2) = 3(x - 1) c. (y - 1) = -3(x - 2) d. (y + 2) = 3(x + 1) Thanks in advance! :-)

OpenStudy (anonymous):

The slope \[m = \frac{y1-y0}{x1-x0} = \frac{5-2}{2-1} = 3\], so, \(y - 2 = 3(x-1)\)

Nnesha (nnesha):

\[\huge\rm \frac{ y_2 - y_1}{ x_2 -x_1 }\] use two order pair where x and y values are \[\huge\rm (x_1 ,y_1)(x_2 ,y_2)\]

Nnesha (nnesha):

\[\huge\rm y -y_1 = m(x - x_1)\] point slope form where m is slope and replace y_1 and x_1 by one of order pair doesn't matter which one you pick

OpenStudy (anonymous):

@Nnesha @M4thM1nd thank you, both! :-) The answer is indeed D.

Nnesha (nnesha):

how is it D ?

OpenStudy (anonymous):

@Nnesha whoops! I mean B!!! I have no idea why I typed that before LOL